# Efficiently serving large multimedia models using EPD Disaggregation

### v1

@@ 3.3.3 Optimized Resource Allocation
The system dynamically determines configurations such as batch sizes, scheduling strategies, and parallelization approaches for each stage of the pipeline. This ensures that critical performance metrics are kept near optimal, based on the workload samples collected during deployment. These workload samples are gathered in real time from a recent time window. If the optimal performance metrics vary significantly in subsequent time windows, the system may adjust the configuration of the EPD pipeline accordingly.
We leverage a black-box optimization algorithm to select configurations, which minimizes GPU usage and maximizes a critical performance metric. For online serving scenarios, we define goodput as the key performance metric. Goodput represents the maximum request rate at which 90% of requests meet the TTFT and TPOT requirements. In offline scenarios, where TTFT and TPOT constraints are less critical, we use E2ETP as the performance metric. By the nature of performance metrics, the optimized config will minimize pipeline inefficiencies (e.g., idle time) and ensure allocated resources are utilized efficiently.
The optimization problem is formulated as:
@@
where $\mathcal{X}$ is the search space for system configurations, including parallelization configs $\mathbf{p}$, max batch size configs $\mathbf{b}$, and scheduling configs $\mathbf{s}$. These variables are vectors, with each element representing the configuration of an individual instance. Here, an instance can handle requests independently, encompassing the (sub-)workers for tensor parallelism and pipeline parallelism. Since instances within a stage process different requests in parallel, we term it data parallelism. Note that $\mathbf{p},\mathbf{b},\mathbf{s}$ can have variable lengths, as the number of instances is itself a configurable parameter. For $i$-th instance, we denote its stage as $\text{Stage}_{i}\in\{E,P,D\}$.
  • Parallelization: Let $\mathbf{p}$ denote the vector of parallel configurations for all instances. For the $i$-th instance, if it is a prefill or decoding instance, then its configuration $p_{i}$ includes: $p^{\text{TP}}_{i}$, the number of GPUs used for tensor parallelism; and $p^{\text{PP}}_{i}$, the number of GPUs used for pipeline parallelism. If it is an encoding instance, considering intra-request parallism (IRP) does not require communication, which is better than TP, we only use IRP. Therefore, we overload the symbol $p^{\text{TP}}_{i}=p^{\text{IRP}}_{i}$ to denote the number of GPUs used for IRP. If the cost per GPU is a constant $c$, then the total cost is: $cost(\mathbf{p})=c\sum_{p_{i}\in\mathbf{p}}(p^{\text{TP}}_{i}\times p^{\text{PP}}_{i})$.
  • Scheduling involves two main questions: First, to which workers or queues should the requests be assigned? Second, how should the order of requests be determined within a queue? We adapt strategies from the DistServe framework [24]. In the encoding stage, when a request arrives, it is assigned and pushed to an instance queue. In the prefill and decode stages, a global queue is used, and each engine pulls from this queue when available. Possible assignment strategies in the encoding stage include Round-Robin or Least-Loaded First for assigning requests. Once a request is assigned to a worker’s queue, we can apply ordering strategies like first-come-first-serve or shortest-job-first, or more complex strategies that prioritize requests based on their Service Level Objectives (SLOs). For simplicity, we constrain that all instances within the same stage share the same scheduling strategy.
In the objective of the problem 6, $\beta$ is a weight that controls the trade-off between performance and cost reduction. In a black-box setting, the performance function $f(\cdot)$ is unknown and must be evaluated through simulations to obtain the performance metric values. To solve the optimization problem, existing methods such as random search, evolutionary algorithm, Bayesian optimization [3], and parametric iterated local search [8] can be employed to find the optimal configuration. We adopt random search to find a near-optimal config for its simplicity.
4 Implementation
@@ 5 Experiments
In this section, we analyze and compare the performance of the proposed EPD disaggregation method against the baselines.
Baseline Methods: We compared our proposed method EPD against two popular baselines: DistServe [24] and vLLM [10]. The DistServe baseline implements the prefill-decode (PD) disaggregation approach, where the prefill and encoding are executed on one set of GPUs, while the decode phase is disaggregated on a separate GPUs. In contrast, the vLLM baseline adopts a monolithic architecture, where all three stages runs on the same set of GPUs. In addition, we also consider an ablation of our method EPD without the IRP feature, named as EPD-IRP.
Since DistServe was originally designed for LLMs, we extended its framework to support LMMs. This required extending the request handling infrastructure to process multimodal data, modifying the block manager to accommodate multimodal tokens, and changing the execution engine from SwiftTransformer to vLLM, which natively supports multiple LLMs. The change to vLLM also ensured consistency in execution engine across all baselines, eliminating potential confounding performance differences arising from variations due to the execution engine.
Models: We utilized three LMMs in our analysis: MiniCPM-V 2.6 [21], InternVL 2.0 8B, and InternVL 2.0 26B [5]. These models are renowned for their advanced capabilities in processing and understanding multimodal data.
MiniCPM-V 2.6 integrates a SigLip-400M vision encoder, comprising 400 million parameters, with a Qwen2-7B language model, containing 7.6 billion parameters, culminating in a total of 8 billion parameters. It excel in tasks involving single-image, multi-image, and video comprehension, even surpassing GPT-4V in these domains [21].
InternVL 2.0 8B features an InternViT-300M-448px vision encoder with 300 million parameters, paired with an internlm2_5-7b-chat language model comprising 7.7 billion parameters, totaling 8 billion parameters. Similarly, InternVL 2.0 26B combines an InternViT-6B-448px-V1-5 vision encoder with 6 billion parameters and an internlm2-chat-20b language model containing 20 billion parameters, resulting in 26 billion parameters. These configurations result in models with 8 billion and 26 billion parameters, respectively.
Datasets: To analyze its performance across diverse scenarios, we utilized custom synthetic workload for all experiments. This approach offered the flexibility to systematically vary parameters such as prompt length, number of images per request, image resolution, output length, and sampling settings. This controlled experimental setup allowed us to assess the system’s robustness and adaptability under different workload configurations, ensuring a comprehensive evaluation of its efficiency and scalability. The length of input tokens in the prompt (excluding multimodal tokens) was fixed to 22. The length of output tokens were fixed to 10 in end-to-end experiments and 1 in the first token generation experiments.
MiniCPMv 2.6 8B InternVL 8B InternVL 26B
#I/R TTFT TPOT TTFT TPOT TTFT TPOT
2 1.40 0.04 1.20 0.05 3.50 0.07
4 2.60 0.04 2.40 0.06 7.05 0.08
6 3.90 0.06 3.55 0.09 11.00 0.95
8 5.10 0.06 5.00 0.18 15.00 0.15
Table 1: Thresholds for TTFT and TPOT (in seconds) used for SLO computation across various models and number of images per request (#I/R) in the experiments.
Evaluation Metrics: We evaluate the system based on runtime performance and memory consumption. The runtime performance is assessed using the TTFT and TPOT metrics for online settings, and E2ETP for offline settings. These metrics are defined as follows:
  • TTFT: The time elapsed from the creation of a request in the Request Generator until the first token is received by the Decoder. This metric captures the responsiveness of the system to incoming requests.
  • TPOT: The time elapsed from the arrival of the first token at the Decoder to the arrival of the last token at the Output Consumer, divided by the number of output tokens. This metric reflects the per-token latency for generating outputs.
  • E2ETP: The end-to-end throughput of the system, measured in terms of requests completed per second. This metric evaluates the system’s ability to handle high request volumes in offline settings.
  • SLO Attainment: The percentage of requests that meet predefined Service Level Objectives (SLOs), such as TTFT and TPOT thresholds. This metric assesses the system’s ability to deliver consistent performance under operational constraints.
  • Goodput: The rate of successfully completed requests that meet SLO criteria, measured in requests per second. This metric emphasizes the system’s effective processing capacity, excluding failed or non-compliant tasks.
For memory benchmarking experiments, we analyze the baselines by evaluating the benefits of available free memory. Additional free memory can facilitate higher batch sizes, accommodate more images per request, or enable larger key-value (KV) cache sizes. We demonstrate the advantages of our proposed approach using these memory-related metrics, highlighting its ability to enhance system scalability and performance.
Implementation Details We conducted our experiments using a cluster of 8 NVIDIA A100 GPUs (82GB), which provided the computational resources necessary for evaluating the proposed system. Each server was equipped with 128 CPUs and 1TB of RAM. The CUDA version was 12.2. Flash attention-2 was used for the attention implementation. We use FP16 precision for all experiments.
To ensure a fair comparison, we standardized key performance-affecting settings of the inference engine across all baselines. Specifically, these include a block size of 16; a maximum of 2048 blocks per request; context tokens capped at 49,152, and decoding tokens at 81,920 per batch. The scheduling policy for all stages was set to First-Come-First-Served (FCFS). Further, to allow enough resources for memory-heavy multimdodal requests to execute, KV cache GPU utililization was set to 50%, and the maximum number of multimedia data of 32 was imposed per prompt. The size of multimodal cache was fixed to 3000 across all models, and the vLLM inference engine was run in eager mode. Finally, for the vLLM inference engine, we used the version 0.6.1.post1 that represents a stable version for multimodal inference.
@@
Figure 4: SLO attainment for end-to-end inference across various models and image counts per request. Column (a) presents results for the MiniCPMv 2.6 model with 2, 4, 6, and 8 images per request (from top to bottom). Column (b) and (c) show results for InternVL2-8B and InternVL2-26B, respectively. The image resolution is fixed at 4K. As observed, EPD significantly outperforms all baselines across all configurations. EPD-IRP means that IRP is disabled in EPD system.
@@ 5.1 SLO Attainment for End-to-End Generation
In this experiment, we evaluate the end-to-end SLO performance of the proposed method (EPD), an ablation of our method without IRP (EPD-IRP) and existing state-of-the-art baselines, DistServe and vLLM. The setup involves 8 NVIDIA A100 GPUs serving 100 multimodal requests, arriving via a Poisson process with a fixed rate ($\lambda$) representing the number of requests per second. We tested three different models—MiniCPMv 2.6, InternVL2-8B, and InternVL2-26B—under varying image counts per request All images were fixed at a 4K resolution of 4032x3024 pixels. The SLO criteria for each configuration, shown in Tab. 1, establish latency upper bounds for both TTFT and TPOT, ensuring timely token generation and consistent decoding performance.
The results for MiniCPMv 2.6 (left), InternVL2-8B (middle), and InternVL2-26B (right) are presented in Fig. 4. Each row corresponds to a different image count per request, specifically 2, 4, 6, and 8, listed from top to bottom. The X-axis represents the per-GPU request rate (requests per second, req/s), while the Y-axis shows the SLO attainment percentage, defined as the proportion of requests meeting both TTFT and TPOT thresholds. The black dotted line indicates the 90% SLO attainment threshold, and the corresponding request rate represents the goodput for 90% SLO attainment.
EPD-IRP demonstrates significantly higher SLO attainment compared to other baselines. At lower request rates, EPD-IRP achieves over 90% SLO attainment in most scenarios. This performance advantage is attributed to EPD-IRP’s ability to parallelize the computationally intensive image encoding step across multiple GPUs, mitigating the delay due to computationally-heavy encoding step. The second-best performer, EPD, achieves an SLO attainment close to 50% in many cases. However, the prior work, DistServe and vLLM, perform the worst, often attaining SLO percentages below 10%. These baselines suffer from interference between encoding and prefill stages, resulting in consistently poor SLO attainment. On the other hand, EPD can mitigate this interference due to disaggregation of encoding and prefill, however it lacks the parallelization ability of EPD-IRP.
As the image count per request increases (from top to bottom rows), the SLO attainment for all baselines declines, particularly at higher request rates. This decline reflects the growing computational demands of processing multiple high-resolution images within a single request. In the most demanding scenario, with 16 images per request (bottom-most row), SLO attainment drops significantly for all baselines. However, EPD-IRP continues to maintain reasonable SLO attainment, especially at lower request rates, highlighting its robustness under heavy workloads.
Finally, as we move from left to right across the plots, the change in model leads to progressively worse performance for all baselines. The InternVL models are prefill-heavy, with prefill stages taking substantially longer due to the generation of a higher number of image tokens. As a result, TTFT suffers from additional queuing delays at the prefill stage, especially at higher request rates. This effect becomes more pronounced when transitioning from the 8B (middle) to the 26B (right) model. In contrast, the MiniCPMv (left) model is optimized to generate fewer image tokens, making it encoding-heavy rather than prefill-heavy. Consequently, it avoids the queuing delays observed in the InternVL models and exhibits better latency characteristics overall.
@@
Figure 5: Distribution of TTFT (y-axis) for various #Images/ Request (x-axis) for the MiniCPMv 2.6 (left), InternVL2-8B (middle), and InternVL2-26B (right) models. Requests arrive following a Poisson distribution with a fixed request rate ($\lambda$). Note that, when comparing TTFT, vLLM is equivalent to DistServe, hence it is skipped. EPD-IRP means that IRP is disabled in EPD system.
@@ 5.2 First Token Generation Latency
In this experiment, we analyze the TTFT performance of the proposed method (EPD), an ablation of our method without IRP (EPD-IRP) and an existing state-of-the-art baseline, DistServe. Unlike prior analyses that focused on end-to-end inference, this experiment isolates the prefill phase of the computation, as the generation of the first token represents a critical latency bottleneck due to large number of prefill tokens in multimodel workloads. Note that in this setup the DistServe baseline is equivalent to vLLM since the decoding phase is not involved.
Fig. 5 shows the results for the three models: MiniCPMv 2.6, InternVL2-8B, and InternVL2-26B, shown in the left, middle, and right plots, respectively. Each subplot presents box plots of TTFT distributions, with the x-axis representing the number of images per request and the y-axis showing the TTFT. The results are generated using 100 requests issued at a request rate of 0.25 for the MiniCPMv 2.6 and 0.08 for the InternVL2-8B and InternVL2-26B models. We limit the number of images per request (x-axis) on the basis of the respective model’s capacity.
As shown in the figure, EPD-IRP significantly outperforms the two baselines, EPD and DistServe, across all three models. Specifically, the TTFT can be upto 71.9%, 32.8%, and 44.9% lower than the DistServe baseline for the MiniCPMv 2.6, InternVL2-8B, and InternVL2-26B models, respectively. This performance advantage stems from its intra-request parallelization, which enables the multimedia data within each request to be encoded in parallel, whereas the baselines process this data sequentially.
Notably, EPD also demonstrates a slight performance improvement over DistServe, highlighting the independent benefits of disaggregating the encoding and prefill stages, even without full parallelization. In this case, the TTFT can be upto 9.8%, 4.9%, and 1.0% lower for the MiniCPMv 2.6, InternVL2-8B, and InternVL2-26B models, respectively.
Further, we also observe that the margin of improvement becomes increasingly pronounced as the number of images per request grows. These results highlight the ability of EPD-IRP to efficiently handle the computational demands of multimodal workloads, especially as the number of input images and overall complexity increase.
@@ 5.3 Memory Savings through Stage Disaggregation
This section demonstrates the significant memory savings achieved by disaggregating the encoding and prefill stages in the EPD system. In existing systems, such as DistServe and vLLM, the encoding and prefill stages are co-located on the same GPUs, requiring each GPU to load both the MME and the LLM. This co-location results in higher memory consumption, as both models must be loaded simultaneously on the same GPUs.
In contrast, the proposed disaggregated system optimizes memory usage by assigning the encoding stage to GPUs that load only the MME, while the prefill stage is handled by separate GPUs that load only the LLM. This separation significantly reduces memory requirements for both stages.
The memory savings achieved through this disaggregation are substantial. Specifically, the encoding-stage GPUs experience memory reductions of approximately 95%, 96.2%, and 78.3% for the MiniCPMv 2.6, InternVL2-8B, and InternVL2-26B models, respectively. On the other hand, the prefill-stage GPUs achieve memory savings of about 5%, 3.7%, and 21.6% for the same models. In practice, the memory savings could be as high as 15x lower as encoding stage do not need to load the KV cache which is only required for prefill. These reductions in memory usage enable the EPD system to allocate resources more efficiently, allowing for the support of additional tasks and models, as demonstrated in the subsequent experiments.
EPD Supports a Higher Number of Images per Request: Tab. 2 compares the maximum number of images per request supported by the disaggregated (EPD) and aggregated (DistServe or vLLM) systems across three multimodal models: MiniCPMv, InternVL2-8B, and InternVL2-26B. The experiment was conducted at three different image resolutions, with a fixed batch size of 1 and 80% of the available memory allocated for KV cache.
For the MiniCPMv model, the EPD approach demonstrates a significant advantage in handling a higher number of images per request. At a resolution of 313x234, EPD supports up to 490 images per request, compared to just 77 for the aggregated DistServe system, resulting in a more than 6x improvement. At the 787x444 resolution, EPD can handle 165 images per request, which is a 6.3x increase over DistServe’s limit of 26. Even at the highest resolution (4032x3024), EPD supports 49 images per request, while DistServe is limited to only 7, demonstrating a 7x improvement.
For the InternVL2-26B model, the improvements are equally striking. EPD supports up to 10x more images per request at both the 313x234 and 4032x3024 resolutions, and approximately 4x more at the 787x444 resolution, compared to DistServe.
Interestingly, for the InternVL2-8B model, both EPD and DistServe support the same number of images per request across all resolutions (19). This is because the InternVL2-8B model model uses a relatively simple MME that is not memory-intensive, and the limit of 19 images is constrained by the maximum context length of the LLM rather than memory usage of the MME.
@@
Table 2: Comparison of maximum supported number of images/request. Batch size is fixed to 1. KV cache is allocated as 80% of the available free memory. Experiment done on A800 GPU with 82GB memory. Higher (italicized) is better.
EPD Supports Higher Batch Sizes for Encoding and Prefill:
Tab. 3 compares the maximum supported batch sizes for the encoding (E) and prefill (P) stages across three multimodal models: MiniCPMv, InternVL2-8B, and InternVL2-26B, at three different image resolutions. The number of images per request is fixed at 10, and each model uses a different patching strategy, leading to varying numbers of image patches. The last three columns show the maximum batch size that can be supported for each configuration without encountering out-of-memory (OOM) errors. For the EPD approach, batch sizes for encoding and prefill are listed separately, as these processes run independently on separate GPUs. In contrast, for the DistServe baseline, encoding and prefill share the same batch size since both stages are executed sequentially on a single GPU.
The EPD approach significantly outperforms the DistServe baseline in terms of supported batch sizes. For instance, with the InternVL2-26B model at a resolution of 787x444, EPD supports a batch size of 22 for encoding, while the DistServe baseline can only handle a batch size of 1, representing a 22x improvement. On the other hand, with the MiniCPMv2.6 model at a resolution of 787x444, EPD supports a batch size of 29 vs 2 for the DistServe baseline, representing a 14.5x improvement for prefill.
@@
Model Image Reso. #Patch E, P E P

### v2

@@ 3.2.3 Optimized Resource Allocation
The EPD system requires tuning various configurations, including batch sizes, scheduling strategies, and parallelization methods for each pipeline stage. To determine the optimal setup, we collect historical workload samples and apply a black-box optimizer.
Let $f(\cdot)$ represent the system’s performance metric (e.g., goodput, see Section 4). Since $f(\cdot)$ is treated as a black-box function with an unknown internal mechanism, we rely on a simulator—extended from DistServe [21]—to evaluate performance metrics efficiently. The objective is to maximize performance while minimizing GPU usage, which inherently reduces pipeline inefficiencies (e.g., idle time) and improves resource utilization. Formally, we solve
@@
Here, $\mathcal{X}$ is the search space for system configs, including parallelization configs $\mathbf{p}$, max batch size configs $\mathbf{b}$, and scheduling configs $\mathbf{s}$ (See Appendix A for details). We use Bayesian optimization [3] to solve Problem 1.
3.2.4 Dynamic Role Switching
The configuration optimizer described in the previous section can determine the optimal settings for a given workload. However, in an online environment, workload characteristics can change dynamically, requiring adjustments to the configuration. Re-initializing the entire system from scratch in response to these changes can be both difficult and costly. For instance, there may be ongoing requests in the encoding, context, or decoding phases with precomputed KV and/or VE caches stored in memory. A naive re-initialization would not only involve booting time but also force these requests to restart from the beginning, potentially causing a cascading impact on SLO metrics for subsequent requests.
To overcome these challenges, we introduce dynamic role switching, which enables any instance in the E, P, or D stages to switch roles to any other stage (E, P, or D) with minimal overhead. At a high level, dynamic role switching continuously monitors the system’s queuing statistics across all stages and reallocates workers to stages experiencing higher demand. When a decision is made to transfer an instance from a source stage $S$ to a destination stage $T$, the migration process occurs in three key steps:
  • Offload: The instance in the $S$ stage stops accepting new requests and redistributes its queued tasks to sibling instances in the same stage.
  • Migration: The instance is reconfigured to meet the requirements of the $T$ stage. This may involve switching both the model and cache type. For example, if the E stage is involved, the instance may switch from an LLM to an MME model and from a KV cache to an MM cache.
  • Onload: The migrated instance resumes processing new requests, helping to alleviate the queuing bottleneck at the $T$ stage.
This process typically takes less than 0.7 seconds. The duration is longer for migrations involving the E stage due to model and cache changes, but significantly shorter when switching between P and D stages, as both the LLM and KV cache can be reused.
@@ 4 Experiments
In this section, we analyze and compare the performance of the proposed EPD disaggregation method against various baselines. We start with an end-to-end generation performance analysis in Sec. 4.1, followed by an examination of first token latency in Sec. 4.2. We then evaluate the memory savings achieved by EPD in Sec. 4.3. Finally, an ablation analysis of key system components is presented in Sec. 4.4. Additional experiments are detailed in the Appendix, including 1) a throughput comparison in offline settings (Appendix C); 2) an extension of our framework to Neural Processing Units (NPUs) demonstrating a 10% performance improvement (LABEL:sec:appendix-npu), and 3) additional SLO and memory experiments (Appendix D).
@@
Figure 5: SLO attainment for end-to-end inference across various models and image counts per request. Columns (a), (b), and (c) correspond to the MiniCPM-V 2.6, InternVL2-8B, and InternVL2-26B, respectively. The top and bottom rows are with 2 and 4 images per request, respectively. As observed, EPD significantly outperforms all baselines across all configurations.
Baselines: We compared the proposed EPD method against two popular baselines: DistServe [21] and vLLM [9]. The DistServe baseline implements the prefill-decode (PD) disaggregation approach, where the prefill and encoding are executed on one set of GPUs, while the decode phase is disaggregated on separate GPUs. Since DistServe was originally designed for LLMs, we extended it to support LMMs by enabling multimodal data processing and modifying the block manager to accommodate multimodal tokens. The vLLM baseline adopts a monolithic architecture, where all three stages run on the same set of GPUs.
Models: We utilized three LMMs in our analysis: MiniCPM-V 2.6 [19], InternVL2-8B, and InternVL2-26B [5]. These LMMs are renowned for their advanced capabilities in processing and understanding multimodal data. Detailed descriptions of the LMMs and their sizes can be found in Appendix B.2.
Datasets: To evaluate performance across diverse scenarios, we conducted experiments using two datasets. The first was a synthetic workload, that allowed setting custom parameters such as prompt length, number of images per request, image resolution, output length, and sampling settings. Unless otherwise specified in our experiments, the input prompt length was set to 22 tokens. The second dataset was NextQA [18], a benchmark video question-answering dataset. Unlike synthetic workload, NextQA features human-annotated questions and answers, providing a more realistic reflection of real-world video request distributions.
Evaluation Metrics: We evaluate the system based on runtime performance and memory consumption. The related performance metrics are listed below.
  • TTFT: The time from the submission of a request to the system until the first token is received by the user.
  • TPOT: The average time interval between each output token (excluding the first token).
  • SLO Attainment: The percentage of requests that meet predefined SLOs, such as TTFT and TPOT requirements.
  • Goodput: The hightest request rate at which 90% or more SLO attainment is achieved.
For memory benchmarking experiments, we analyze the baselines by evaluating the benefits of available free memory. Additional free memory can facilitate higher batch sizes, accommodate more images per request, or enable larger key-value (KV) cache sizes. (See section 4.3)
@@ 4.1 SLO Attainment for End-to-End Generation
In this experiment, we evaluate the goodput of EPD and the baselines in an online setting, where 100 multimodal requests arrive following a Poisson process with rate $\lambda$. The length of output tokens is fixed to 10. We test requests with 2 and 4 images, each at a resolution of $4032\times 3024$. The details of the corresponding SLO criteria are discussed in Sec. B.3. The results for three models (columns) are presented in Fig. 5. The X-axis represents the per-GPU request rate ($\lambda$/#GPUs), and the Y-axis indicates the percentage of requests meeting both TTFT and TPOT requirements, with a 90% threshold denoted by a black dotted line.
@@
Figure 6: Distribution of TTFT (Y-axis) for various #Images/ Request (X-axis) for the MiniCPM-V 2.6 (left), InternVL2-8B (middle), and InternVL2-26B (right)
EPD outperforms all baselines, achieving over 90% SLO attainment at lower request rates. This is due to its ability to parallelize the computationally intensive image encoding step across multiple GPUs. DistServe and vLLM often maintain less than 10% attainment due to interference. Meanwhile, comparing the first and second rows for varying image counts per request (2 and 4), we observe that while more images per request increase workload, EPD maintains reasonable performance, whereas baseline SLO attainment drops significantly. Lastly, InternVL, being prefill-heavy, experiences queuing delays due to the higher number of image tokens, particularly at higher request rates. This effect exacerbates performance degradation from the 8B to the 26B model. In contrast, MiniCPM-V, optimized to generate fewer image tokens, avoids these delays and achieves better overall latency. Further, we provide more results with 6 and 8 images in Sec. D.1.
Next, we repeat the experiment using the non-synthetic video question-answering dataset, NextQA [18]. For this, we randomly sampled 100 examples, with input text token lengths ranging from 4 to 21 (average: 11.42) and output token lengths ranging from 1 to 7 (average: 2.75). Each video request was represented by 8 uniformly sampled frames, and we used MiniCPM-V 2.6 in the experiment, adhering to the SLO criteria of TTFT = 5.60 and TPOT = 0.06. As illustrated in Fig. 7, EPD is the only framework achieving 90% SLO attainment at low request rates, demonstrating its superior effectiveness in handling diverse, real-world-like workloads compared to DistServe and vLLM.
Figure 7: SLO attainment for end-to-end inference on the NextQA dataset. As seen, EPD outperforms other baselines by a significant margin.
@@ 4.2 First Token Generation Latency
Multimodal requests often place a heavy load on both the E and P phases, making the first token generation a key latency bottleneck. Therefore, in this experiment, we analyze the first token latency for various baselines. The results presented in Fig. 6 shows box plots of TTFT distributions for three different models. Note that, as the decoding phase is excluded, the vLLM baseline is equivalent to DistServe and is thus skipped. Requests are generated according to a Poisson distribution with a fixed request rate $\lambda$. Specifically, $\lambda=0.25$ for the MiniCPM-V 2.6, and $\lambda=0.08$ for both the InternVL2-8B and InternVL2-26B. Thanks to intra-request parallelization, EPD significantly outperforms both baselines, EPD and DistServe. Specifically, TTFT is reduced by up to 71.9%, 32.8%, and 44.9% compared to the DistServe baseline for the MiniCPM-V 2.6, InternVL2-8B, and InternVL2-26B models, respectively. The performance gain grows with the number of images per request, demonstrating EPD’s efficiency in handling heavy multimodal workloads.
@@ 4.3 Memory Savings through Stage Disaggregation
In this section we analyze the memory savings achieved by disaggregating the encoding and prefill stages. The E workers can save memory as they do not require LLM weights or the KV cache. Analyzing only the weight size indicates memory reduction of approximately 95%, 96.2%, and 78.3% for the MiniCPM-V 2.6, InternVL2-8B, and InternVL2-26B models. Similarly, for the P workers, memory saving of about 5%, 3.7%, and 21.6% are achieved, respectively. In practice, since KV cache is also not required at E workers, the memory saving can be even higher (93.3% saving, i.e., 15$\times$ lower, according to profiling). These reductions in memory usage enable the EPD system to support higher numbers of images per request and batch sizes (discussed in the following experiments) and larger KV cache sizes (discussed in Appendix D.2).
@@
Table 1: Comparison of maximum supported number of images/ request. Batch size is fixed to 1. KV cache is allocated as 80% of the available free memory. Higher (italicized) is better.
EPD Supports a Higher Number of Images per Request: We compare the maximum number of images per request supported by the disaggregated (EPD) and aggregated (DistServe or vLLM) systems across three multimodal models: MiniCPMv, InternVL2-8B, and InternVL2-26B. The experiment was conducted at three different image resolutions, with a fixed batch size of 1 and with 80% of the available memory allocated for KV cache. As shown in Tab. 1, EPD handles more images per request than DistServe, for example, at 4032$\times$3024 resolution, 7$\times$ more for InternVL2-26B and 10$\times$ more for InternVL2-8B. For InternVL2-8B, the limit of 19 images per request is due to its maximum context length. Without this constraint, a larger number of images per request could be supported.
EPD Supports Higher Batch Sizes: We compare maximum supported batch sizes for E and P stages across different settings of image resolution and models in Tab. 2. As seen, EPD significantly outperforms DistServe for both E and P batch sizes. For example, with InternVL2-26B at 787$\times$444 resolution, EPD supports a batch size of 22 for encoding vs. DistServe’s 1, achieving 22$\times$ improvement. Similarly, for MiniCPM-V 2.6 at 787$\times$444, EPD achieves a batch size of 29 vs. DistServe’s 2 for prefill, a 14.5$\times$ improvement.
@@
Model Image Reso. #Patch (E, P) E P